OpenStudy (anonymous):

Line AB contains points A (0, 0) and B (2, 2). Line CD contains points C (3, 1) and D (5, 3). Lines AB and CD are parallel , OR perpendicular , OR neither

OpenStudy (anonymous):

find the slope of line AB with the formula y2-y1/x2-x1. do the same for line CD. if the slope is the same, they are parallel. if they are the opposite reciprocals, they are perpendicular.

OpenStudy (anonymous):

@d-jay @TechnoSoul @matt101 @TashaMeni7 @HelpBlahBlahBlah

OpenStudy (anonymous):

i got parallel

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

Line AB contains points A (8, −4) and B (1, −5). The slope of line AB is −7 negative 1 over 7 1 over 7 7 @surjithayer

OpenStudy (anonymous):

The equation of line AB is y = 5x + 1. Write an equation of a line parallel to line AB in slope-intercept form that contains point (4, 5). y = 5x − 15 y = 5x + 15 y = 1 over 5x + 21 over 5 y = 1 over 5x − 29 over 5

OpenStudy (anonymous):

find same way

OpenStudy (anonymous):

im lost wats the answers i have 3 mins on my online homewolrk

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@urkhada

OpenStudy (anonymous):

eq. of line parallel to y=5x+1 is y=mx+c if (4,5) lies on it,find c

OpenStudy (anonymous):

correction y=5x+c